OpenStudy (anonymous):

what will be the ratio between 3.3*10^-18 and 9.9*10^-18?

OpenStudy (mathstudent55):

\( \dfrac{3.3 \times 10^{-18}}{9.9 \times 10^{-18}} \) \(= \dfrac{3.3 }{9.9 } \times \dfrac{10^{-18}}{ 10^{-18}} \) \(= \dfrac{3}{9} \times \dfrac{1.1 }{1.1 } \times \dfrac{10^{-18}}{ 10^{-18}} \) \(=\dfrac{1}{3} \times \dfrac{3}{3} \times \dfrac{1.1 }{1.1 } \times \dfrac{10^{-18}}{ 10^{-18}} \) \(=\dfrac{1}{3} \times \dfrac{\cancel{3}}{\cancel{3}} \times \dfrac{\cancel{1.1} }{\cancel{1.1} } \times \dfrac{\cancel{10^{-18}}}{ \cancel{10^{-18}}} \)

OpenStudy (anonymous):

Thanks!

OpenStudy (mathstudent55):

wlcm